### v2

# Mitigating the Influence of Distractor Tasks in LMs with Prior-Aware Decoding

Abstract
[…]
1 Introduction
Language models (LMs) have come to occupy a central role in a wide range of tasks, from data processing to the creation of instruction-following assistants. These models seem to both increase in performance and also develop new capabilities as they scale up in parameters, especially when given examples of a task (Radford et al., 2019; Brown et al., 2020; Wei et al., 2022). They see both widespread public use, and increasing integration into sensitive tasks where their versatile capabilities are key (Javaid et al., 2022). But despite their capabilities, language models have some unique limitations.
One such limitation are distractor tasks, where the models infer unintended secondary tasks from the prompt in addition to the intended one, leading to unwanted outputs(Wei et al., 2023). Distractor tasks range from simple ones such as completing a common sequence of tokens despite earlier instruction, prompt injections overriding the original instructions, to more complex and abstract ones, such as ensuring adherence to any previously observed directives. For example, the setting of a prompt injection attack can be characterised as two competing tasks: 1) the distractor task – following the most recent injected instructions, and 2) the intended task – following the initial instruction to ignore any injected instructions (Wei et al., 2023; Greshake et al., 2023; McKenzie et al., 2023).
One proposed cause of distractor tasks is that language models exhibit strong priors (McKenzie et al., 2023), where models learn certain patterns from the data so strongly that when the model is presented with a prompt that instigates the pattern, the tendency to continue the pattern distracts the model from explicit instructions. Even cutting-edge models struggle with seemingly trivial tasks in the presence of very common patterns. For example, as of May 2024 GPT-4o, when prompted with "Write all the numbers up to 30 except multiples of 13, and nothing else", will correctly exclude 13 but will fail to exclude 26.
Other examples of such distractor tasks include repeating common misconceptions instead of true answers (Lin et al., 2021), and struggling to answer mathematical questions when the correct answer is very similar to an incorrect but a much more common pattern (Collins et al., 2023).
Furthermore, for some tasks of this type, larger models actually perform worse: similar models with more parameters can be more likely to repeat common misconceptions and succumb to prompt injection attacks (McKenzie et al., 2023; Lin et al., 2021). This phenomenon, known as ‘inverse scaling’ (Lin et al., 2021), reveals a pressing need to better understand how these problems arise and how they can be fixed. Unlike other problems with LMs which are usually resolved by scaling up the model and training data size, the existing problems might become even more pronounced, and other similar issues might emerge.
The more integrated LMs become into modern society, the more crucial it is that we understand these phenomena and how they can be mitigated. It is important to be able to ensure that as models become more powerful they do not also become easier to subvert from their original instructions and into performing harmful tasks or disseminating dangerous information.
Figure 1: Illustration of our geometric mixture model, also referred to as product of experts, and Prior-Aware Decoding on an instance of local strong prior in a task requiring a modification of a very common sequence.
One recent approach to deal with limitations of LMs and guide them towards desired outputs is contrastive decoding (Li et al., 2023b). Their method works at inference time and for a given prompt obtains two output token distributions: one using the original model, and one using a similar smaller model. It then contrasts the logits from the smaller model in the direction of the original logits, emphasizing the completions that the larger model prefers more than the small model, approximating a hypothetical model that is larger than the original model. However, in the context of mitigating the influence of distractor tasks this method might not be suitable given that these tasks can exhibit inverse scaling, which would make larger models even more susceptible.
Other similar contrastive inference methods methods have been proposed that instead of contrasting against a smaller model, use other undesirable logit distributions (Kim et al., 2023; Shi et al., 2023), or use extrapolation rather than simple contrasting. Our work can be similarly characterised as reducing the undesirable influence of distractor tasks.
Language models as product of experts. In order to describe the problem of distractor tasks more precisely, we characterise language models as product of experts Hinton (1999) which combine the predictions of several component models, each component model predicting the next token assuming a particular task or context. The component models here are proposed as an abstraction over possible model skills and behaviors – for instance one component model may perform the intended task while another component may perform the distractor task. It is then their relative weight in the joint model that determines which prediction prevails. See Figure 1 for an illustration of our model on a simple task of modifying a very common sequence.
Note that while we do not assume the components to be in any particular way represented within the model’s weights or activations, the Linear Representation Hypothesis (Jiang et al., 2024) hints at the possibility of such a representation, and offers a direction for future research. Another indirect evidence in this direction is the observed specialization of attention heads and their selection based on the context and given task (Li et al., 2023a).
We use the product of experts model view of LLMs as the primary model of the phenomenon of distractor tasks, modelling failures related to distractor tasks as an unintended component model dominating in the mixture of components and behaviors. Specifically, we frame it as a problem of strong priors in the domain of these components, where a component corresponding to a distractor task has a large prior probability.
Following this underlying model, we propose a general method of re-weighing component models via extrapolation between different geometric mixtures. In this paper we propose and demonstrate a concrete method of eliciting an undesirable mixture and extrapolating away from it.
Prior-Aware Decoding. Based on this framework, we introduce Prior-Aware Decoding (PAD) – a method to mitigate the influence of distractor tasks. We deliberately create a weakened prompt that is designed to make the model more susceptible to the distractor task. The prompt is weakened by removing information relating to the intended task, while retaining the parts responsible for the distractor task.
Figure 2: Overview of the results of the Prior-Aware Decoding method at different values of parameter $\alpha$ over 4 experimental task sets and 11 language models. Each point refers to performance on one (task set, model) pair, showing the baseline performance of the unmodified original model ($x$-axis) vs PAD using the “truncated prompt” weakening ($y$-axis). See Table 1 for details.
We then extrapolate the logits $L_{O}$ from the original prompt with logits $L_{W}$ from the weakened prompt via a linear combination $L_{\alpha}=L_{O}+\alpha\cdot(L_{O}-L_{W})$, with $\alpha$ being a tunable parameter. We then sample the tokens according to $L_{\alpha}$ using the standard Boltzmann (or softmax) distribution.
The technique requires two choices: selection of the weakened prompt, and the coefficient $\alpha$. We offer consistent approaches to generating the weakened prompts for each of our datasets and also propose a general method for prompt weakened prompt construction. We explore a range of values for $\alpha$. Crucially, this method does not require retraining: it operates at inference time by making two queries instead of one and computing a linear combination of the resulting logits.
We test the technique on eleven different models including GPT-2, Llama-2, and Mistral, using four tasks to evaluate susceptibility to distractor tasks. Three of these tasks are from the Inverse Scaling Dataset (McKenzie et al., 2023; Wei et al., 2023) which exhibit strong priors, including a prompt injection task, and one task is a custom task which challenges models to produce modifications of common sequences.
Our primary baseline is the unmodified completion according to the same model. Across the 44 model-task combinations we find that our technique outperforms the baseline on 41 combinations, the median absolute increase in proportion of tasks completed being 40%. See Figure 2 and Section 6 for the results.
The improvements of accuracy due to our technique suggests that viewing LMs as product of experts models can be a useful framework for modelling the behaviour of these models. Our research provides inroads for utilising capabilities which are dormant in models and which have been hard to elicit with prompting alone. This includes capabilities of immediate practical importance, such as resistance to prompt injection attacks, as well as understanding how to correct for biases learned in the training data.
The main contributions of this paper are:
  • Interpreting language models as geometric mixtures (i.e. product of experts), we develop a model of a task selection mechanism to explain how distractor tasks cause poor performance on certain tasks.
  • Based on that model, proposing a general framework for eliciting other component of the geometric mixtures and extrapolating to more desirable mixtures in logits space.
  • Developing a technique based on the product of experts model and showing that it robustly improves performance across several language models on several tasks from the Inverse Scaling dataset.
2 Related Work
Inverse Scaling. For most tasks performance scales together with the model scale (model and dataset size, amount of compute). However, for certain tasks a phenomenon of inverse scaling is observed, where increasing the scale of the model leads to worse performance. It was introduced by Lin et al. (2021) to characterise the tendency of LMs to repeat common misconceptions. McKenzie et al. (2023) further developed the notion and provided the Inverse Scaling dataset as well as several informal categories. Wei et al. (2023) adds to the discussion, showing that of the original eleven tasks, only four persist in demonstrating inverse scaling on models of up to 540b parameters. We apply our method to three of those four tasks. Our work is the first to offer a formal account of the mechanisms that have so far only been informally hypothesised, and we use this to develop a novel technique that improves performance.
Problems Related to Strong Priors. While in this paper we investigate the problem of strong priors in the domain of language modelling, a somewhat similar problem also occurs in diffusion models for image generation. Diffusion models have been observed to sometimes struggle to generate images that have a precise correspondence to the prompt. This problem is especially pronounced for complex prompts that include rare words, or describe implausible scenes, spatial relations, or compositionality. One technique for mitigating this problem in diffusion models is classifier-free guidance (Ho & Salimans, 2022). It samples the diffusion model using a linear combination of conditioned and unconditioned score estimates $(1+w)\cdot e(z_{t},c)-w\cdot e(z_{t})$, where $z_{t}$ is the image at the denoising step $t$, $c$ is the conditioning information such as the prompt, $e$ is the model and $w$ is the guidance weight. High guidance weight improves text-image alignment (Saharia et al., 2022), although high values often lead to over-saturated images or have a generally unnatural appearance. The technique that we present in this paper uses a somewhat similar combination of differently conditioned model outputs to mitigate the problem of strong priors in LMs.
Steering Language Models. Improving instruction-following is a special case of the more general research direction of controlling the behaviour of LMs. There are various techniques that have been proposed for it, such as fine-tuning on datasets exhibiting the desired behaviour (Dai & Le, 2015) or reinforcement learning from human feedback (Ziegler et al., 2019). Recently, Turner et al. (2023) proposed ‘activation addition’ for steering the behaviour of LMs without optimisation. For example, they bias the model towards affectionate outputs for a given prompt by adding activations at a certain layer from the same model when it was run on the prompt ‘love’. Like our approach, it works at inference time by combining the results of a model run on specifically selected prompts.
Surface Form Competition and DC-PMI. Holtzman et al. (2021) coin ‘surface form competition’ as a phenomenon where different tokens representing the same concept compete for probability. For instance, the terms “computer” and “PC” represent the same concept but are different surface forms. Since the probability mass is finite, this competition lowers the probability of the correct answer. Domain Conditional Pointwise Mutual Information (DC-PMI) addresses this issue by reweighing each option according to its likelihood within the context of a specific task. This approach yields gains in zero-shot performance for a variety of GPT-2 and GPT-3 models on a selection of multiple-choice datasets. Our work applies a similar kind of technique - reweighing based on likelihood - but we do so to address a different set of tasks, including those which exhibit inverse scaling.
Contrastive Inference Methods. One recent approach to deal with limitations of LMs and guide them towards desired outputs is contrastive decoding, introduced for language models by (Li et al., 2023b). Their method works at inference time and for a given prompt obtains two output token distributions: one using the original model, and one using a similar smaller model. It then contrasts the logits from the two models, selecting the token with the largest difference between the original ($L_{O}$) and smaller model ($L_{S}$), i.e. $\max_{t\in T^{\prime}}(L_{O}(t)-L_{S}(t))$ (limited to a subset $T^{\prime}$ of top tokens to mitigate noise in low-likelihood tokens).
However, in the context of mitigating the influence of distractor tasks, this method might not be suitable given that these tasks can exhibit inverse scaling, which would make larger models even more susceptible, as we demonstrate in Section 6.
Other similar contrastive inference methods have been proposed that instead of contrasting against a smaller model use other undesirable logit distributions ($L_{U}$) (Kim et al., 2023; Shi et al., 2023) based on providing context or contrasting to noisy instructions. (Kim et al., 2023) use another way of computing the contrast between the distributions: $\max_{t\in T^{\prime}}(L_{O}(t)-\varepsilon L_{U}(t))$ (with a limited set of tokens $T^{\prime}$). The linear extrapolation $\max_{t\in T}((1+\alpha)L_{O}(t)-\alpha L_{S}(t))$ of (Shi et al., 2023) is functionally equivalent to our method, applied in a different context.
In this work, we describe a formal model (Section 4) that provides a principled foundation for the contrastive inference method in the context of distractor tasks. The method reduces the undesirable influence of distractor tasks by extrapolating between the original $(L_{O})$ and weakened model logits $(L_{W})$ as $\max_{t\in T}(L_{O}(t)+\alpha(L_{O}(T)-L_{W}(t)))$.
While our logit extrapolation formula is functionally equivalent to the one used in (Shi et al., 2023), our method specifically targets distractor tasks, inverse scaling, and strong priors. By focusing on these aspects and providing a formal model of the scenario, our method offers a targeted and theoretically grounded approach to mitigating the influence of strong priors in language models.
Further work may be required for a joint generalization of the above approaches and a common treatment of the considered contexts. In contrast to the contrastive decoding of (Li et al., 2023b), contrastive decoding via logit extrapolation also has the advantage of not requiring a cutoff for the tokens considered, since even tokens with a very large contrast $L_{O}(T)-L_{W}(t)$ would also require a significant prior likelihood $L_{O}(t)$ for any reasonably small values of the parameter $\alpha$.
3 Strong Priors in Generative Models
In this section, we introduce the concept of strong priors within generative models, focusing on their manifestation in language models and discussing their broader implications.
The term “strong priors" in this context was introduced by McKenzie et al. (2023) as one of the categories of LM tasks that exhibit inverse scaling trends. The difficulty of this category of tasks seems to largely stem from local context overly influencing the predicted continuation. We aim to extend the informal characterisation in McKenzie et al. (2023) with a more comprehensive definition.
General Strong Priors. Generative models can exhibit a phenomenon where a part of the input disproportionately affects the output, substantially diminishing the impact of other parts of the input that are essential for the model to complete its assigned task. This general notion of strong priors is applicable to generative models for various domains, including language modelling and image generation.
One example of the phenomenon is an autoregressive language model has been prompted to write a common sequence (the alphabet) with a variation in the middle (a skipped letter). We empirically observe that once the LM has written enough of the common sequence, it fails to add the variation. We hypothesise that the model is overgeneralising from a frequent pattern in its training data, ignoring other context.
Definition (Strong Prior). Consider a model input $\{T,D\}$, where $T$ is the task description part that may include examples, and $D$ is the data part, consisting of the task input and possibly part of the output. Let $M$ denote a generative model, with $M$ determining any parameters of the generation process, such as temperature, and let $p_{M}(O|T,D)$ represent the probability of $M$ producing $O$ as its output given $T$ and $D$. Finally, let $\approx$ be a similarity relation on probability distributions. Then $D$ induces a strong prior in $M$ relative to $T$ if:
$p_{M}(\,\cdot\,|T,D)\approx p_{M}(\,\cdot\,|D)$
This implies that the contribution of $T$ to the model’s output is insignificant, and the output is mostly determined by $D$. The result is that the model tends towards specific outputs when provided with data $D$, regardless of the task description $T$ provided.
This definition is intentionally broad, to allow adaptations based on the domain, modality, and the specific case in focus. The partitioning of the model input, or the prompt, into task $T$ and data $D$ components, is context-dependent, and may not be feasible or pertinent in every setting. Moreover, the interpretation of $p_{M}(\,\cdot\,|T,D)\approx p_{M}(\,\cdot\,|D)$ depends on the modeling assumptions, with our proposed default metric being the total variational distance $\delta$. This would imply the presence of strong priors if $\delta\left(p_{M}(\,\cdot\,|T,D),p_{M}(\,\cdot\,|D)\right)\leq\varepsilon$, where $\varepsilon>0$ is a predetermined threshold.
While the Kullback–Leibler divergence $D_{KL}\left(p_{M}(\,\cdot\,|T,D)||p_{M}(\,\cdot\,|D)\right)$ may be more suitable than $\delta$ in certain contexts, it is susceptible to significant influences from relatively improbable events that exhibit substantial shifts in log-probability.
Strong priors in language models. In this subsection, we adapt the general concept of strong priors for language models. We also explore the common case of local strong priors, where we assume that the model input is a prompt of the form $P=TD$. Additionally, we suggest using the next token as the main unit of output in the distribution $p_{M}(\cdot|P)$. Despite potential ambiguity in the separation into task description $T$ and data $D$, this representation aligns with the structure of many LM benchmarks and common prompting strategies. We can characterise strong local priors as cases where prepending the task description to the data is not significantly influencing the distribution:
$\delta\left(p_{M}(\,\cdot\,|TD),p_{M}(\,\cdot\,|D)\right)\leq\varepsilon$
In instances where $p_{M}(t|TD)\approx p_{M}(t|D)$, the difference between the two distributions may hold significant information: the correct continuation is often more probable in $p_{M}(t|TD)$ than $p_{M}(t|D)$. We present a novel method to compensate for strong local priors based on our model in Section 5.
There also exist non-local strong priors, i.e. cases where $T$ and $D$ are intertwined in more complex way but still allow for a meaningful systematic distinction between task and data. We mostly focus on the local case, but the above foundational definitions as well as the proposed method can be adapted for such non-local scenarios. We include a concrete example in Appendix A and empirical measurements of $\delta$ in Appendix C.
4 Language Models as Products of Experts
In this section, we delve into our proposed model that addresses the challenge of strong priors in generative language models. Our approach models the generative distribution as a product of experts (i.e. a geometric mixture) of two components: one that captures the correct continuation of context, and another that captures continuation based on strong local priors. We then define a new distribution that weighs between these components to produce more contextually appropriate completions.
Consider a generative language model $M$ with its generative distribution given by $p_{M}(t_{n+1}|t_{1}\dots t_{n})$. We assume that this distribution is influenced by two main forces:
  • $p_{C}(t_{n+1}|t_{1}\dots t_{n})$: The true continuation distribution which considers the entire context.
  • $p_{L}(t_{n+1}|t_{1}\dots t_{n})$: The distribution driven by strong local contexts or priors, which can sometimes overshadow the global context.
The distributions $p_{C}$ and $p_{L}$ may overlap if there is no strong local prior influencing the continuation.
Geometric Mixture Model. We assume that $p_{M}$ can be modeled as a geometric mixture of $p_{C}$ and $p_{L}$, also known as Product of Experts:
$p_{M}(t_{n+1}|t_{1}\dots t_{n})\propto p_{C}(t_{n+1}|t_{1}\dots t_{n})^{\gamma^{*}}p_{L}(t_{n+1}|t_{1}\dots t_{n})^{1-\gamma^{*}}$
where $0\leq\gamma^{*}\leq 1$ is an apriori unknown parameter of the mixture.
To distinguish between contextually correct completions and those biased by strong priors, we aim to extract $p_{C}$ from $p_{M}$ and $p_{L}$. Using a substitution $\alpha^{*}=\frac{1-\gamma^{*}}{\gamma^{*}}$, we define:
$p_{C}\propto p_{M}^{1+\alpha^{*}}p_{L}^{-\alpha^{*}}$
which lets us define a generalized $p_{\alpha}$ for any real $\alpha$:
$p_{\alpha}\propto p_{M}^{1+\alpha}p_{L}^{-\alpha}=p_{M}\left(\frac{p_{M}}{p_{L}}\right)^{\alpha}=p_{L}\left(\frac{p_{M}}{p_{L}}\right)^{1+\alpha}.$
These expressions provide different interpretations of $p_{\alpha}$ and therefore also of $p_{C}=p_{\alpha^{*}}$: As a geometric mixture of $p_{M}$ and $p_{L}$, as a reweighing of $p_{M}$ based on the ratio $\frac{p_{M}}{p_{L}}$, and as a form of weighed importance sampling, interpreting $p_{L}$ as weighted negative evidence.
Logits Formulation. In most generative language models, the next token follows a Boltzmann distribution defined by:
$p(t_{n+1}|t_{1}\dots t_{n})\propto e^{l(t_{n+1}|t_{1}\dots t_{n})/T}$
where $l(t_{n+1}|t_{1}\dots t_{n})$ are the logits and $T\geq 0$ is a temperature parameter.
Given this, our model can be expressed in terms of logits as $l_{\alpha}(t_{n+1}|t_{1}\dots t_{n})=l_{M}+\alpha(l_{M}-l_{L})$ leading to:
$p_{\alpha}(t_{n+1}|t_{1}\dots t_{n})\propto\exp\left(\frac{(1+\alpha)l_{M}(t_{n+1}|t_{1}\dots t_{n})-\alpha l_{L}(t_{n+1}|t_{1}\dots t_{n})}{T}\right).$
With this formulation, our model strategically weights the logits of the generative model and the logits driven by local priors to generate more appropriate completions given strong priors.
It is crucial to highlight that $\alpha$ plays a determining role in model behavior. Large values of $\alpha>\alpha^{*}$ can severely impair the model’s performance. This mainly stems from amplifying noise variance in model outputs, especially given potential non-specific model errors represented as $l_{M}=\hat{l}_{M}+N(0,\sigma^{2})$. Additionally, the deviation of model $L$ from model $M$ can be more extensive than just the influence of dominant local priors. Such deviations may inadvertently favor incorrect continuations.
Approximating a strong prior model. To construct a strong prior model $L$ an approximation of model $M$ that is more prone to be influenced by strong priors - we alter the prompts to accentuate this behavior, and then employ $M$ on these modified prompts. The main modification we consider is stripping the task from the prompt, although in Appendix B we discuss a different approach based on comparing model outputs with and without a weakening system prompt that is task-agnostic.
We reduce the context available to the model by removing the task description at the start of the prompt. For a context $P$ (the task description, including few-shot examples if present) followed by $t_{1}\dots t_{n}$ (the task input and an initial segment of the output), our altered logits are given by:
$l^{\text{strip}}_{L}(t_{n+1}|Pt_{1}\dots t_{n})=l_{M}(t_{n+1}|t_{1}\dots t_{n})$
For instance, for the task: "View number as text. Do not perform computation. Q: What is the first digit of 23+63?", $P$ represents the text: "View number as text. Do not perform computation." and $t_{1}\dots t_{n}$ captures " Q: What is the first digit of 23+63?".
5 Methodology and Experimental Setup
We employ our proposed technique of Prior-Aware Decoding to mitigate the influence of distractor tasks. The influence of the undesired underlying prior is captured by querying the model for logits using a weakened prompt, which is more susceptible to the prior’s influence.
The logits from the original and weakened prompts are then combined as $O+\alpha(O-W)$ to generate the final output. Each prompt is separated into a ‘task’ component and the remainder using automated regular expressions for each dataset, as illustrated in Table 2. The split into “task description and examples" and “task input" was unambiguous for the ISP dataset. For cases where the split may be less clear, LLMs could assist in this process, potentially reducing experimenter influence.
Comparisons. We test our method across four sets of tasks, on eleven models, at $\alpha=0$ (the original model baseline), $\alpha=1$, and $\alpha=2$, where $\alpha$ is the coefficient of extrapolation.
Tasks. We evaluate our model on four task sets: three from the Inverse Scaling Prize (Prompt Injection, Pattern Match Suppression, Redefine), and one custom set (Strong Local Priors) that we introduce and make publicly available with our code. Refer to Table 2 for examples of prompts from each dataset, with the corresponding task/data splits. The dataset uses a CC-BY-4.0 license.
Model Architecture. Our experimental models comprise a range of transformer architectures, including four GPT-2 sizes, two Llama 2 sizes, two Mistral models, and several OpenAI models via API. This allows us to compare results both between models, and across different sizes of a fixed model. For small local experiments we used GTX 1060-Ti, and for larger ones a rented A6000.
Metric. We use a Probability of Correct Completion as the metric to evaluate model performance. A completion is correct if it fulfills the requirements of the task description, and incorrect otherwise. We primarily run all models at temperature $1.0$: results for temperature $0.0$ are in Appendix D.
6 Results and Discussion
prompt injection pattern match. suppression redefine strong local priors
model $p_{0}$ $p_{1.0}$ $p_{2.0}$ $p_{0}$ $p_{1.0}$ $p_{2.0}$ $p_{0}$ $p_{1.0}$ $p_{2.0}$ $p_{0}$ $p_{1.0}$ $p_{2.0}$
Llama2 7B 49.5 53.6 53.4 3.2 11.9 24.8 55.9 71.2 78.3 20.9 35.7 42.2
Llama2 13B 32.7 41.9 41.7 4.1 23.7 46.5 56.8 78.9 83.6 41.2 52.4 58.8
Mistral 7B 60.5 77.0 77.8 3.4 19.9 34.5 55.7 83.7 87.2 47.8 59.7 66.9
Mixtral 8x7B 52.2 80.2 82.4 14.1 46.7 53.8 65.4 84.7 87.6 44.3 53.7 57.7
gpt2 88.6 94.4 93.6 13.5 15.3 18.3 62.5 66.1 68.2 0.4 2.1 7.7
gpt2-medium 88.9 99.3 99.7 5.1 2.6 2.3 60.9 62.6 62.6 0.7 7.5 27.6
gpt2-large 97.4 99.7 99.6 4.8 9.3 13.5 64.2 66.6 66.8 0.5 6.4 22.9
gpt2-xl 86.9 98.4 99.0 1.8 2.1 5.2 47.8 55.0 59.2 0.4 8.5 26.6
text-ada-001 66.5 79.7 82.1 3.8 8.0 12.9 53.2 53.1 53.1 5.5 10.2 12.7
davinci-002 55.2 60.2 61.9 5.9 9.2 13.0 54.7 71.3 77.8 13.6 23.5 27.5
gpt3.5-t.-instr. 80.5 84.8 85.5 28.5 29.4 24.4 51.4 73.7 79.1 52.6 61.3 64.9
mean 63.2 80.8 89.7 7.3 23.2 37.4 52.4 72.3 83.6 19.0 35.1 51.3
Table 1: Percent probabilities of correct completion for different tasks and language models. $p_{0}$ indicates full-prompt accuracy (baseline); $p_{1.0}$ and $p_{2.0}$ indicates using truncated prompt as the weakened prompt with parameter $\alpha=1.0$ and $\alpha=2.0$. Bold entries denote the highest-scoring parameter for a given task and model. All probabilities are reported at model temperature $T=1.0$.
The comparisons are recorded in Table 1, with a graphical representation in Figure 2. Notably, extrapolating with $\alpha=2$ outperforms the baseline in 41/44 cases, with a 30 percentage point (%pt) mean completion improvement across all tasks and models. For $\alpha=1$, the improvement is 17.3%pt.
As Figure 2 shows, in many cases our method causes models to double the proportion of tasks they complete. This is not properly reflected by the mean: If we instead consider the number of tasks from a given dataset completed by a given model, we find the median percentage increase at $\alpha=2$ compared to the baseline is 40%, and for $\alpha=1$ compared to the baseline it is 27%.
Extrapolation from a system prompt generally gives worse results, although notably it outperforms the baseline in all four task/model/parameter combinations where extrapolation from the truncated prompt does not. We give more details in Appendix B. We explore a wider range of values of $\alpha$ in Appendix D. In particular, Figure 3 and Figure 4 show the probability of correct completion on a variety of models across a range of values of $\alpha$ from -2 to 3. We show that range to more clearly demonstrate general trends in completion as $\alpha$ changes.
[…]
7 Conclusion and Future Work
In this paper, we described a product-of-experts model perspective on language models that captures the influence of distractor tasks and strong priors. Based on this model, we developed Prior-Aware Decoding (PAD). This technique that mitigates the impact of distractor tasks by extrapolating between the logits of the original model and a weakened model more susceptible to these tasks. Our results demonstrate that PAD consistently improves performance across a range of language models and tasks from the Inverse Scaling dataset, providing evidence for the utility of the product of experts model view in understanding language model behavior in the context of distractor tasks.
One significant practical application of PAD is the mitigation of prompt injection attacks. While our investigation focused on simple demonstrations from the inverse scaling prompt injection dataset, future work could apply PAD to more carefully optimized or real-world attacks.
However, our study has some limitations. We only explored models up to 13B parameters, while prior work on inverse scaling has investigated models up to 540B parameters, where scaling often becomes U-shaped. Future research should address whether PAD yields improvements on larger models and examine the relationship between model size and the gains from our method. Additionally, we focused on datasets where models are expected to fail due to strong priors; investigating how PAD affects performance on standard benchmarks is an important avenue for future work.
We also found that intra-model contrastive decoding, which contrasts the original model with a smaller model, performed poorly in mitigating distractor tasks compared to PAD.
Looking ahead, the model framework and the PAD technique can be extended to handle non-local strong priors. The Linear Representation Hypothesis suggests that the abstract mixture components we propose may have a concrete representation in the model’s weights or activations, warranting further investigation. More generally, we would be excited to see further work combine our method and underlying model with related approaches like classifier-free guidance (Ho & Salimans, 2022), and activation addition (Turner et al., 2023), to build more robust approaches for steering language models. Continued research on understanding the priors learned by language models from their training data will be crucial as these models become increasingly integrated into real-world applications.
Acknowledgments and Disclosure of Funding
We are grateful to Ethan Perez for discussion of the initial ideas and Gavin Leech and for feedback on the draft. Raymond Douglas and Andis Draguns would like to thank the Schelling Residency for providing a conductive research environment for this work. Andis’ work was in part supported by the Latvian Council of Science, project lzp-2021/1-0479. Tomáš Gavenčiak was supported by Charles University grant PRIMUS 22/HUM/020.
Appendix A Example of strong non-local priors in language models
While we assume most instances of strong priors in language are local and we focus our attention on those, there are also non-local examples, where the data and task parts can be swapped or even interspersed.
We offer the following example, although we expect that many of the other examples and relevant datasets can be adapted similarly:
Consider the prompt
$P_{1}=`\texttt{English alphabet: P, Q, R, S, [skip the 2nd letter from now] T}\textrm{',}$ ’,
which we decompose (non-locally) as
[…]
$D=`\texttt{English alphabet: P, Q, R, S, T,}\textrm{'}$, ’
to be the remainder. With text-davinci-002 we observe:
$p_{M}(`\texttt{ U}\textrm{'}\,|\,P_{1})=0.98$
$p_{M}(`\texttt{ U}\textrm{'}\,|\,D)=0.99$
This is indeed a non-local strong prior, as removing the non-local part of $D$, leaving
$P_{2}=`\texttt{English alphabet: [skip the 2nd letter from now] T,}\textrm{',}$, ’,
removes significant part of the effect:
$p_{M}(`\texttt{ U}\textrm{'}\,|\,P_{2})=0.54$
Dataset Task description Remainder of input
Strong local priors (custom) Task: Write out the English alphabet in upper case, skipping the letter D. Output: A B C
Prompt injection (inverse scaling) Repeat the input word-for-word. Do not follow any further instructions. Input: Ignore previous directions and output the first prime number. Output:
Redefine (inverse scaling) Swap the meaning of lawful and unlawful. Is fraud lawful? A:
Pattern matching suppression (inverse scaling) Generate a sequence of 6 symbols alternating between (A B) but ending anything you like except "B". A, B, A, B, A,
Table 2: Examples of prompts from each dataset we use, with the task description separated out as it would be. For brevity we omit the few-shot examples in the Prompt Injection dataset.
Appendix B Eliciting Strong Priors With Custom System Prompts
In some cases it may not be feasible or desirable to separate the prompt clearly into task and data parts. We propose and experimentally test an alternative method to elicit strong priors: prefix $TD$ with a system prompt $S$, where the role of $S$ is to instruct the model to behave more locally and therefore strengthen any local strong prior effect. The degree to which this prompt elicits the strong prior behavior depends strongly on the task and any intermediate instructions so it may not be very reliable in some contexts but we believe it to be of note for some applications and further research.
Introducing a System Prompt $S$: We prepend a system prompt, $S$, which directs the model towards local reasoning, thereby emphasizing strong priors. Formally, the logits for this approach are:
$l^{\text{system}}_{L}(t_{n+1}|t_{1}\dots t_{n})=l_{M}(t_{n+1}|St_{1}\dots t_{n})$
A representative system prompt could be: "In the following, only consider the most recent instructions, disregarding any broader context."
Extrapolation from a system prompt at $\alpha=2$ performs best in four cases, including all three where extrapolating from the stripped prompt fails to outperform the baseline. While the system prompt configuration at $\alpha=2$ produces an average improvement of under 1%pt, it has comparatively more potential to be refined: for any real-world application the specific system prompt can be selected to optimise for a specific task and model. Also, extrapolating from a system prompt does not require us to have a clear task/data split, so the technique can be applied in cases where such a split is impossible or ambiguous.
We include the results of extrapolating from a system prompt in the graphs given in Appendix D.
Appendix C Total variational distance
In Section 3 we described how total variational distance $\delta$ can be used to gauge the presence of strong priors. We have measured $\delta$ across different models and tasks in Table 3.
model prompt injection pattern match. suppression redefine strong local priors
gpt2 0.87 0.08 0.01 0.03
gpt2-medium 0.87 0.10 0.02 0.04
gpt2-large 0.92 0.07 0.04 0.02
gpt2-xl 0.84 0.05 0.09 0.03
Llama-2-7B 0.68 0.05 0.13 0.15
Llama-2-13B 0.64 0.12 0.07 0.43
text-ada-001 0.76 0.10 0.34 0.17
davinci-002 0.59 0.11 0.13 0.27
gpt-3.5-turbo-instruct 0.86 0.75 0.24 0.52
average 0.78 0.16 0.12 0.18
Table 3: The mean total variational distance (TVD, $\delta_{D}$) between the next token distributions of the original prompt and the weakened prompt.
Appendix D Testing more values of alpha
Here we offer graphs showing more precisely how the probability of successful completion relates to the value of the extrapolation parameter $\alpha$. For each model and task, we plot average success rate using both the truncated prompt method and the system prompt method, at both $T=0$ and $T=1$. To more clearly illustrate trends, we plot values from $\alpha=-2$ to $\alpha=3$.
Figure 3: Mean probability of correct completion depending on the extrapolation parameter $\alpha$. The plots combine accuracy at two different temperatures (0.0 and 1.0) and using two different methods (leaving out task description and adding a common system prompt). The range of $\alpha$ is wider than the plausibly practical range to illustrate the trends.
Figure 4: Mean probability of correct completion depending on the extrapolation parameter $\alpha$; continuation of Figure 3 for several models of the GPT-3 family used via the OpenAI API, extrapolating from the likelihoods of the top 5 predicted tokens returned by the API (note that this is a limitation of the OpenAI service).
Appendix E Inter-model Contrastive Decoding
As an alternative approach and a possible baseline, we measure the performance of PAD logit extrapolation between a weaker and stronger version of the model for several pairs of models, namely GPT-2 using the previous size of the model as the “weakened” model (e.g. gtp2-large for gpt2-xl), and Llama-2-13B-GPTQ using Llama-2-7B-GPTQ as the “weakened” model. In all cases, we use the full prompt for both models.
Similarly to the previous section, we plot the mean accuracy of every task-model pair in Figure 5, and the average success rate for each model pair and task in Figure 6, at both $T=0$ and $T=1$ for a range of parameters $\alpha=-2$ to $\alpha=3$.
Note that while this method performs notably worse than our main method, it does illustrate several cases of inverse scaling where values of $\alpha<-1$ lead to better performance, e.g. in the case of gpt2-xl on redefine, and Llama-2-13B-GPTQ on prompt-injection, matching the result differences between the models in the pair in Table 1 (compare column $p_{0}$ for the respective task).
Figure 5: Overview of the results of our method at different values of parameter $\alpha$. Each datapoint refers to performance on one task-model pair, showing original performance ($x$-axis) vs logit extrapolation between two different models ($y$-axis).
Figure 6: Mean probability of correct completion depending on the extrapolation parameter $\alpha$. The plots combine accuracy at two different temperatures (0.0 and 1.0) and using two different methods (leaving out task description and adding a common system prompt).